The ratio of the capacities of two similar tanks are 64:125. If 4.8 kg of paint is needed to paint the smaller tank, how many kilograms of paint will needed for the bigger tank?

a. 9.4

b. 7.5

c. 5.7

d. 8.6
Answer #1

Let the weight of paint needed for the bigger tank be x kg.

Since, the tanks are similar.

Therefore, \frac{64}{125}=\frac{4.8}{x}

or, x=\frac{125X4.8}{64}=9.4

Answer: Option a